# Sequential Diversification with Provable Guarantees

### v2

@@ 8.2. Experimental setting
Distance function. For the recommendation datasets, given the categorical nature of the data, we use the Jaccard distance as the item-item distance function. In particular, if $\mathcal{C}_{i}$ denotes the set of categories of item $i$, the Jaccard distance $d(i,j)$ is defined as $d(i,j)=1-{\lvert\mathcal{C}_{i}\cap\mathcal{C}_{j}\rvert}/{\lvert\mathcal{C}_{i}\cup\mathcal{C}_{j}\rvert}$. For the information-retrieval datasets, with the provided document feature vectors, we use the $1-$cosine distance as the document-document distance function.
[…]
@@ 8.3. Proposed methods and baselines
EXPLORE (Coppolillo et al., 2024). Details on the EXPLORE algorithm and its adaptation to our setting are provided in the appendix.
@@
[…]
$\textsc{B}\tau\textsc{I}$. We refer to the $\textsc{B}\tau\textsc{I}$ algorithm as B2I, B3I, and B4I when $\tau$ is set to be $2$, $3$, and $4$. Once the first $\tau$-sequence is chosen, we greedily extend the remaining items that maximize the incremental gain of MaxSSD. Notice that B2I in our implementation is the same as the greedy algorithm we propose in Algorithm 3.
[…]
Throughout the experiments, we evaluate B2I, B3I, and B4I on dataset LETOR and LTRC. For the recommendation datasets, we evaluate B2I and heuristics B3I-H and B4I-H on these datasets.

### v3

@@ 9.2. Experimental Setting
Distance function. For the recommendation datasets, given the categorical nature of the data, we use the Jaccard distance as the item-item distance function. In particular, if $\mathcal{C}_{i}$ denotes the set of categories of item $i$, the Jaccard distance $d(i,j)$ is defined as $d(i,j)=1-{\lvert\mathcal{C}_{i}\cap\mathcal{C}_{j}\rvert}/{\lvert\mathcal{C}_{i}\cup\mathcal{C}_{j}\rvert}$. For the information-retrieval datasets, with the provided document feature vectors, we use the 11-cosine distance as the document-document distance function.
Additional metrics. To evaluate the performance of our algorithms in terms of user engagement and satisfaction, we use two metrics: expected discounted cumulative gain ($\mathit{Exp\-DCG}$) and expected serendipity ($\mathit{Exp\-Serendip\-i\-ty}$). Discounted cumulative gain (Järvelin and Kekäläinen, 2002) measures the accumulated relevance of items in a ranking list, discounted by their position. Serendipity (Herlocker et al., 2004) quantifies how relevant and unexpected the recommended items are to the user. We adapt these metrics to account for users’ continuation probabilities. For any user, let $p_{i}$ be the continuation probability of item $i$.
$\mathit{Exp\-DCG}$ is defined as $\sum_{j=1}^{|\mathcal{I}|}\left(\sum_{t=1}^{j}\frac{p_{t}}{log_{2}(t+1)}\right)(1-p_{j+1})\prod_{t=1}^{j}p_{t}$. For serendipity, let $H$ be the set of previously rated items, and $\mathcal{C}(H)$ be the set of categories covered by $H$. $\mathit{Exp\-Serendip\-i\-ty}$ is defined as $\sum_{j=1}^{|\mathcal{I}|}\left(\sum_{t=1}^{j}p_{t}\cdot I(t,u)\right)(1-p_{j+1})\prod_{t=1}^{j}p_{t}$, where $I(t,u)=1$ if $\lvert\mathcal{C}(H\cup\{t\})\rvert>\lvert\mathcal{C}(H)\rvert$ and 0 otherwise. Here, $p_{t}$ assesses the relevance of item $t$, while $I(t,u)$ captures its unexpectedness.
Note that, by definition, the $\mathit{Exp\-Serendip\-i\-ty}$ score for the LETOR and LTRC datasets is 0 as the sequences for each user comprise only rated items. Similarly, the KuaiRec dataset also has a $\mathit{Exp\-Serendip\-i\-ty}$ score of 0, as each user has already covered all categories in their history of ratings. Therefore, we omit the $\mathit{Exp\-Serendip\-i\-ty}$ scores for these datasets.
@@ 9.3.1. Baselines
EXPLORE (Coppolillo et al., 2024). Details on the EXPLORE algorithm and its adaptation to our setting are provided in the full paper.
@@
9.3.2. Our Methods
We provide more details of our methods $\small\textsc{B}\tau\textsc{I}$, $\small\textsc{B}\tau\textsc{I-H}$, $\small\textsc{SB}\tau\textsc{I}$ and $\small\textsc{SB}\tau\textsc{I-H}$.
$\small\textsc{B}\tau\textsc{I}$. We refer to the $\small\textsc{B}\tau\textsc{I}$ algorithm as B2I, B3I, and B4I when $\tau$ is set to be $2$, 33, and $4$. Once the first $\tau$-sequence is chosen, we greedily extend the remaining items that maximize the incremental gain of MaxSSD. Notice that B2I in our implementation is the same as the greedy algorithm we propose in Algorithm 3.
$\small\textsc{B}\tau\textsc{I-H}$. Since $\small\textsc{B}\tau\textsc{I}$ can be computationally expensive when $\tau$ is large, we adopt a greedy approach by selecting the top 100 items that maximize the incremental gain of MaxSSD, thereby forming a candidate set. The best $\tau$ items are exclusively selected within this candidate set. We extend the best $\tau$ items in the same way as in $\small\textsc{B}\tau\textsc{I}$.
$\small\textsc{SB}\tau\textsc{I}$. This is the $\small\textsc{B}\tau\textsc{I}$ algorithm with $O_{\tau}=\operatorname*{arg\,max}_{O_{\tau}\sqsubseteq\pi(U)}\mathcal{S}_{+}(O_{\tau})$. We refer to the $\small\textsc{SB}\tau\textsc{I}$ algorithm as SB3I and SB4I when $\tau$ is set to $3$ and $4$, respectively. Note that when $\tau$ is set to $2$, SB2I is identical to B2I.
$\small\textsc{SB}\tau\textsc{I-H}$. Analogous to $\small\textsc{B}\tau\textsc{I-H}$, this is the heuristic version of $\small\textsc{SB}\tau\textsc{I}$, where we select the best $\tau$ items exclusively from the same candidate set as described for $\small\textsc{B}\tau\textsc{I}$.
Throughout the experiments, we evaluate B2I, B3I, B4I, SB3I and SB4I on dataset LETOR and LTRC. For the recommendation datasets, we evaluate B2I and heuristics B3I-H, B4I-H,SB3I-H and SB4I-H on these datasets.
@@ 9.4. Results and Discussion
9.4.1. Performance on Sequential Sum Diversity.
Table 2. Sequential sum diversity $\mathcal{S}_{+}(O)$ with item continuation probability mapped to $[0.1,0.3]$. The results marked with are obtained using the $\small\textsc{B}\tau\textsc{I-H}$ and $\small\textsc{SB}\tau\textsc{I-H}$ heuristic
Coat KuaiRec Netflix Movielens Yahoo LETOR LTRC
Random 0.039 $\pm$ 0.020 0.022 $\pm$ 0.006 0.063 $\pm$ 0.026 0.059 $\pm$ 0.028 0.016 $\pm$ 0.027 0.008 $\pm$ 0.008 0.019 $\pm$ 0.011
EXPLORE 0.103 $\pm$ 0.041 0.050 $\pm$ 0.000 0.152 $\pm$ 0.026 0.173 $\pm$ 0.013 0.133 $\pm$ 0.044 0.026 $\pm$ 0.021 0.041 $\pm$ 0.018
DUM 0.102 $\pm$ 0.038 0.033 $\pm$ 0.008 0.145 $\pm$ 0.026 0.169 $\pm$ 0.018 0.148 $\pm$ 0.044 0.043 $\pm$ 0.028 0.054 $\pm$ 0.020
MSD 0.095 $\pm$ 0.039 0.042 $\pm$ 0.006 0.161 $\pm$ 0.024 0.183 $\pm$ 0.003 0.148 $\pm$ 0.044 0.039 $\pm$ 0.026 0.051 $\pm$ 0.020
MMR 0.105 $\pm$ 0.041 0.051 $\pm$ 0.010 0.162 $\pm$ 0.024 0.183 $\pm$ 0.003 0.148 $\pm$ 0.044 0.040 $\pm$ 0.026 0.051 $\pm$ 0.019
DPP 0.105 $\pm$ 0.041 0.044 $\pm$ 0.006 0.162 $\pm$ 0.024 0.183 $\pm$ 0.003 0.148 $\pm$ 0.044 0.040 $\pm$ 0.026 0.051 $\pm$ 0.019
B2I 0.109 $\pm$ 0.042 0.058 $\pm$ 0.010 0.163 $\pm$ 0.024 0.183 $\pm$ 0.003 0.148 $\pm$ 0.044 0.043 $\pm$ 0.028 0.055 $\pm$ 0.020
B3I 0.106 $\pm$ 0.038* 0.058 $\pm$ 0.009* 0.160 $\pm$ 0.025* 0.183 $\pm$ 0.003* 0.140 $\pm$ 0.045* 0.042 $\pm$ 0.027 0.054 $\pm$ 0.020
B4I 0.104 $\pm$ 0.037* 0.058 $\pm$ 0.009* 0.159* $\pm$ 0.025* 0.183 $\pm$ 0.003* 0.139 $\pm$ 0.045* 0.042 $\pm$ 0.027 0.054 $\pm$ 0.020
SB3I 0.083 $\pm$ 0.046* 0.058 $\pm$ 0.010* 0.164 $\pm$ 0.023* 0.183 $\pm$ 0.003* 0.152 $\pm$ 0.041* 0.044 $\pm$ 0.028 0.054 $\pm$ 0.198
SB4I 0.084 $\pm$ 0.046* 0.058 $\pm$ 0.010* 0.164 $\pm$ 0.023* 0.183 $\pm$ 0.003* 0.152 $\pm$ 0.041* 0.044 $\pm$ 0.028 0.054 $\pm$ 0.198
Table 3. Sequential sum diversity $\mathcal{S}_{+}(O)$ with item continuation probability mapped to $[0.4,0.6]$. The results marked with are obtained using the $\small\textsc{B}\tau\textsc{I-H}$ and $\small\textsc{SB}\tau\textsc{I-H}$ heuristic.
Methods Coat KuaiRec Netflix Movielens Yahoo LETOR LTRC
Random 0.646 $\pm$ 0.177 0.383 $\pm$ 0.066 0.921 $\pm$ 0.220 0.878 $\pm$ 0.246 0.231 $\pm$ 0.248 0.208 $\pm$ 0.106 0.359 $\pm$ 0.116
EXPLORE 1.276 $\pm$ 0.374 0.681 $\pm$ 0.082 1.861 $\pm$ 0.263 2.142 $\pm$ 0.114 1.684 $\pm$ 0.465 0.440 $\pm$ 0.190 0.597 $\pm$ 0.151
DUM 1.214 $\pm$ 0.348 0.411 $\pm$ 0.086 1.779 $\pm$ 0.241 2.080 $\pm$ 0.113 1.793 $\pm$ 0.466 0.564 $\pm$ 0.184 0.640 $\pm$ 0.134
MSD 1.211 $\pm$ 0.349 0.704 $\pm$ 0.057 1.888 $\pm$ 0.254 2.228 $\pm$ 0.046 1.793 $\pm$ 0.466 0.577 $\pm$ 0.191 0.662 $\pm$ 0.146
MMR 1.282 $\pm$ 0.377 0.725 $\pm$ 0.080 1.937 $\pm$ 0.258 2.232 $\pm$ 0.044 1.793 $\pm$ 0.466 0.570 $\pm$ 0.191 0.644 $\pm$ 0.145
DPP 1.284 $\pm$ 0.377 0.711 $\pm$ 0.073 1.933 $\pm$ 0.258 2.230 $\pm$ 0.046 1.793 $\pm$ 0.466 0.557 $\pm$ 0.181 0.642 $\pm$ 0.138
B2I 1.289 $\pm$ 0.376 0.778 $\pm$ 0.075 1.946 $\pm$ 0.264 2.235 $\pm$ 0.043 1.795 $\pm$ 0.466 0.592 $\pm$ 0.195 0.670 $\pm$ 0.144
B3I 1.184 $\pm$ 0.368* 0.776 $\pm$ 0.072* 1.919 $\pm$ 0.270* 2.234 $\pm$ 0.046* 1.729 $\pm$ 0.495* 0.582 $\pm$ 0.197 0.667 $\pm$ 0.145
B4I 1.184 $\pm$ 0.367* 0.782 $\pm$ 0.073* 1.910 $\pm$ 0.272* 2.233 $\pm$ 0.050* 1.715 $\pm$ 0.496* 0.588 $\pm$ 0.197 0.668 $\pm$ 0.145
SB3I 1.076 $\pm$ 0.420* 0.782 $\pm$ 0.078* 1.948 $\pm$ 0.247* 2.228 $\pm$ 0.055* 1.834 $\pm$ 0.448* 0.597 $\pm$ 0.196 0.673 $\pm$ 0.143
SB4I 1.076 $\pm$ 0.420* 0.786 $\pm$ 0.077* 1.951 $\pm$ 0.249* 2.227 $\pm$ 0.054* 1.834 $\pm$ 0.448* 0.600 $\pm$ 0.199 0.677 $\pm$ 0.144